### v4

Abstract
Determining when people are struggling from video enables a finer-grained understanding of actions and opens opportunities for building intelligent support visual interfaces. In this paper, we present a new dataset with three assembly activities and corresponding performance baselines for the determination of struggle from video. Three real-world problem-solving activities including assembling plumbing pipes (Pipes-Struggle), pitching camping tents (Tent-Struggle) and solving the Tower of Hanoi puzzle (Tower-Struggle) are introduced. Video segments were scored w.r.t. the level of struggle as perceived by annotators using a forced choice 4-point scale. Each video segment was annotated by a single expert annotator in addition to crowd-sourced annotations. The dataset is the first struggle annotation dataset and contains 5.1 hours of video and 725,100 frames from 73 participants in total. We evaluate three decision-making tasks: struggle classification, struggle level regression, and struggle label distribution learning. We provide baseline results for each of the tasks utilising several mainstream deep neural networks, along with an ablation study and visualisation of results. Our work is motivated toward assistive systems that analyze struggle, support users during manual activities and encourage learning, as well as other video understanding competencies.
keywords:
Computer Vision Dataset, Egocentric Video, Struggle Determination
1 Introduction
Figure 1: Struggle and non-struggle examples. upper box Three non-struggle examples from Pipes-Struggle, Tent-Struggle, and Tower-Struggle. Three struggle examples from the same activities are shown in the bottom box. Frames are displayed in chronological order from left to right. Non-struggle examples feature fluid motion without errors, whereas videos showcasing struggling show issues (highlighted in orange). For example: repeatedly placing the item back (top row struggle), assembling errors (middle row struggle), and hesitation towards the next step (bottom row struggle).
Understanding when someone is struggling is an important cognitive competence and a key component of skill acquisition [38]. This capability could enable the development of systems that understand how effectively people perform activities and assist with achieving goals or supporting learning outcomes. The ability to detect struggle and subsequently offer help, is a core developmental ability and exists already in pre-verbal children and even other primates [59].
Motivated by the larger goal of finer-grained visual understanding and opportunities on eye-wear computing to provide real-time assistance, we collected a dataset for struggle determination using indoor and outdoor real-world activities, including three domains: assembling plumbing pipes; pitching a camping tent; and solving the ‘Tower of Hanoi’ puzzle. These three domains were chosen specifically to demonstrate the viability of struggle detection, due to their task structures and assembly requirements. The Tower of Hanoi puzzle presented a strict rule-based framework requiring participants to solve a spatial puzzle. Pipe plumbing uses rigid objects and printed instructions for participants to follow, showcasing their ability to perform an unknown task with guidance. The tent assembly task introduced a free-form assembly process with deformable objects, demanding greater improvisation and manual skill. Compared to other general real daily-living tasks (eg cooking or crafts), our chosen tasks are examples of real-world problem-solving and have well-defined goals (completion of the structures). Our tasks allow participants to demonstrate a variety of complex visuomotor behaviours with multiple, albeit constrained, opportunities for struggle.
Our struggle-determination approach is based on two main assumptions. First, we believe there is sufficient information for a large number of activities to determine if a person is struggling from video, based on low-level motion cues of hand-object interaction and without requiring specific task knowledge. This is supported by behavioural evidence e.g. [59, 60] in people. Second, the same deep learning models and train/test strategies can be used for different tasks as they share common visual cues that help to determine struggle, such as the smoothness of hand motion.
Figure 1 shows sample video frames from our data. We used two levels for annotation: MTurk workers and a domain expert. Annotations used a forced choice, four-level struggle scale so that we can determine if a person is struggling and to what degree. We evaluated several video understanding methods trained on our dataset and provided a variety of baseline results across our proposed struggle determination activities and machine learning tasks.
Our primary contributions are: 1) Surfacing the problem of struggle determination for finer-grained assembly understanding in egocentric videos. 2) Introducing a public dataset for struggle determination, which is annotated via an expert annotator and crowdsourcing. 3) Presenting three benchmark tasks for struggle determination: classification, regression, and label-distributed learning, all using weakly supervised learning with only video-level labels. 4) We provide initial benchmark results of video understanding models on struggle determination thus providing baseline results and demonstrating the essential role of motion information.
2 Related Work
We first review related prior video understanding research, highlighting action recognition, action quality assessment, skill ranking and video anomaly detection. We also summarize related datasets and make comparisons to our struggle determination offering, see Table 6 in the Appendix.
Action Recognition. Action recognition (AR) aims to classify actions from video. Many mainstream deep architectures are initially designed for AR but have been used for other video tasks such as Action Segmentation/Detection [24, 6, 67], Action Anticipation [20], and video-text retrieval [32, 17]. Current AR datasets differ in granularity between coarse-grained actions, in which the background context displays large differences between classes [47, 26, 25, 12], and fine-grained actions in which models need to discriminate between action-motion information such as hand-object interactions [7, 66, 19, 27, 46, 62].
Numerous deep architectures have been developed for AR. Initially, 2D-ConvNets were created to distinguish video actions from 2D spatial features in video frames, typically combined with temporal average pooling and/or optical flow to improve motion understanding. Subsequently, 3D-ConvNets [55, 4, 15, 61, 56, 57, 16] gained popularity due to their ability to model motion in the 3D convolution operation, as well as being able to train using large-scale datasets, such as Kinetics [25] and the proposal of expanding 2D convolutions into 3D [4].
Recently, Transformer-based architectures such as Vision Transformer (ViT) [8], Swin Transformer [33], MLP-Mixer [54], and gMLP [30] have been used increasingly. For AR, a hierarchical architecture has been developed to extract spatial-temporal features efficiently, such as ViViT [1], TimeSformer [3], MViT [14, 28], and Video Transformer Networks (VTN) [37].
Action Quality Assessment. Action quality assessment (AQA) is a regression task predicting action quality scores. Most datasets, such as AQA-7 [41], Olympic Scoring Dataset [42, 44], JIGSAW [18], and MTL-AQA [43], contain sports scenes, where annotations represent the scores given by expert human judges. The typical architecture for AQA consists of 3D convolutional neural networks (C3D) [55] to extract spatial-temporal features, followed by average pooling (C3D-AVG), LSTM [22] (C3D-LSTM), or attention modules to further aggregate information from various video segments and generate regression or multitask predictions [43]. Some methods tackle this task using Label Distribution Learning [51] in which the distribution of assessment scores is predicted. Methods often predict the scores directly using MLPs [65], or the mean/std. dev of a Gaussian Distribution based on Variational Auto-Encoders (VAE) [51, 68].
Skill Determination. Similar to AQA, Skill Determination [40, 39, 9, 10, 21, 29] aims to predict the skill level of the action being performed but does so via ranking. [40, 39] mainly focus on ranking surgery skills using Siamese LSTM-based networks by aggregating kinematic data. [9] investigated ranking skills for general daily activities and introduced the EPIC-Skill2018 dataset. [29] proposes an RNN-based spatial attention model and significantly improves the results in [9]. [10] employs a combination of contrastive learning and attention to discriminate between high- and low-skill actions that require fine-grained discrimination skills. Most recently, [21] proposes a TikTok dance video dataset for ranking dancing performances.
Video Anomaly Action Detection. Video Anomaly Detection aims to detect unintentional [11, 64] or anomalous actions [69, 49, 53, 70]. Whilst fine-grained discrimination may share some similarities to Struggle Detection we differ in two major aspects: persons struggling may not contain anomalies and Anomaly Detection is constructed as a temporal localisation task.
We find that fine-grained video understanding has been explored in a variety of different datasets and tasks (see Table 6 in the Appendix). However, albeit crucial for fine-grained activity and skill understanding, determining the level of struggle of the participant has not been attempted. Previous research efforts motivate us to create three tasks for our dataset: struggle classification, struggle level regression, and struggle label distribution learning. As labels for these tasks do not exist, we annotate struggle for one existing dataset [23, 48] and collect and annotate two new activities. Our three activities were chosen to require complex visuomotor coordination and problem-solving while remaining goal-driven. This promotes similar yet varied occurrences of struggle behaviours and provides enough diversity within and across activities.
3 Struggle Determination Dataset
Here we detail the methodology to record our egocentric videos and collect annotations for the struggle determination dataset. Our struggle dataset is presented under the Non-Commercial Government Licence for public sector information [52].
We initially recorded the behaviour of the participants in three scenarios that require complex visuomotor coordination and problem-solving and where participants may occasionally make errors or otherwise struggle. These scenarios were intended to be similar to what one might encounter in daily life (both indoors and outdoors) and where one might want to receive instructional advice. Finding suitable scenarios for struggle detection presents a balancing act. The scenarios shouldn’t be too easy, as participants might not exhibit struggle behaviours, but they also shouldn’t be overly complex or open-ended to prevent unclear goals and excessive task completion times (we favour activities contained within a few minutes). One alternative option would have been to leverage existing visuomotor datasets, such as EPIC-Kitchen [7], by selecting and annotating videos related to struggle. We have investigated this option and generated Table 6 (see in appendix) summarizing potential datasets. Unfortunately, we encountered two key limitations on most previous datasets: (1) Misaligned annotations: None of the existing datasets offered annotations that specifically targeted or purposefully quantified struggle; (2) Scarcity of repeated goal-directed tasks: existing datasets mostly lack repeated instances of goal-directed activities (e.g., not every participant is cracking eggs or flipping pancakes). These limitations led us to consider collecting new data rather than just relying on the existing ones. An exception is [23, 48], where several subjects repeat the same activity for a complex visuomotor assembly of a camping tent. We annotate this data in our work. Finally, we want real scenarios that are ordinary enough so that naive observers can estimate if a participant is struggling without prior knowledge. Across the three activities, we recorded first-person videos and later supplemented the videos with struggle-determination annotations from human observers.
The scenarios included: assembling plastic plumbing pipes by following one of two diagrams (easy or difficult) while seated indoors; solving a version of the Tower of Hanoi puzzle while seated indoors; and finally assembling a camping tent moving freely outdoors. We respectively name the scenarios as Pipes, Tower, and Tent and the corresponding activities/subsets for struggle determination as Pipes-Struggle, Tower-Struggle and Tent-Struggle. Each scenario includes fine-grained actions that people perform routinely (e.g. arranging, picking up and placing objects, attaching them, etc.). Equipment needed, difficulty and goals vary between scenarios, as does participant performance, with some struggling more or less in completing a scenario.
Table 1: Struggle level descriptions.
Definitely non-struggling The actions are executed with complete ease and confidence, with no indication of any challenge or difficulty encountered throughout the task.
Slightly non-struggling The actions are generally executed smoothly and with confidence, with only occasional minor errors or corrections that indicate some level of challenge or difficulty.
Slightly struggling There are some signs of challenge or difficulty including occasional errors, pauses, dropping off items, or wobbly hands that indicate the participants may not be very experienced in the task at hand but are still making an effort to complete it.
Definitely struggling There are clear, consistent signs of difficulty, such as repeated attempts, unsmooth motions, and significant pauses during the task. These actions indicate that the participants don’t really know what to do and may lack the knowledge or skill necessary to complete the task successfully.
3.1 Dataset Construction
3.1.1 Data Collection
For all three scenarios, university undergraduate and graduate student populations were recruited. All participants gave informed consent and approval for data sharing. We recruited 24 individuals to perform the plumbing task (80% female, mean age 22.6 years), and 20 participants from this group also performed the Tower of Hanoi task. For Struggle-Tent, as stated in [23], 29 individuals (71% female, mean age 23.3 years) performed the tent-pitching scenario. In all activities, participants wore a head-mounted GoPro Hero 5 (1080p at 60Hz).
In the plumbing task, participants assembled home plumbing pipes following a diagram of a double sink layout (see supplementary material). Participants were evenly split between those with an easier and more difficult diagram to follow. To simplify dexterity requirements, the activity was reduced to assembling the layout of the pipes on a table rather than in a real sink.
Videos were captured at $1920\times 1080$ resolution, at 30Hz for Tent-Struggle, and 60Hz for Pipes-Struggle and Tower-Struggle. Struggle determination annotation was accomplished by MTurkers rating 10s video clips from the tasks. We segment videos into 10s blocks. Given the struggle’s inherent ambiguity, we discovered during the early annotation process that it can be difficult for annotators to locate the beginning and end of the struggle in untrimmed video recordings. Therefore we chose to use a weakly labelling approach instead. Inspired by Experimental Psychology research (e.g. [59], [60]) that shows pre-verbal children and even chimpanzees can determine struggle actions within 10s, our clips were created by uniformly splitting the videos into 10-second segments from beginning to end (e.g. 0-9s, 10-19s, etc.). We also resized videos to $456\times 256$ at 30Hz. This ensured segments had a sufficient resolution for human observers while maintaining a resolution suitable for deep learning models. Our annotation process thus focuses on evaluating struggle within isolated 10-second clips. This means that we assess and label the presence of struggle solely within each clip, without considering the broader context of the full video recording. Using the EPIC-Tent action label annotations [23], we chose to focus on video segments from five actions (assemble support, insert stake, insert support, insert support tab, and place guyline) where our expert annotator judged they had the most obvious visual cues of struggle in a series of pilot studies.
3.1.2 Annotating Struggle
Since the original videos are trimmed uniformly into 10-second segments we focus on ‘low-level’ struggling cues which we believe to be the most transferable between participants and scenarios. Struggle markers may include shaky hands, pausing actions, dropping items, or action repetition. While there is a variety of ways one might ask an observer to estimate struggle, we chose to ask observers to rate video segments using a four-point scale varying from definitely non-struggle to definitely struggle. Table 1 lists the descriptions for each struggle level. The value of a finer, four-point gradation of struggle supports the longer-term goal of assistive systems. For example, the support provided to a person who is ‘slightly’ struggling may consist of a few basic instructions or reminders. However, a person determined to be definitely struggling may require more detailed guidance or even recommendations to stop the task to prevent injury.
This scoring approach allows naive observers to indicate the degree of struggle/non-struggle. Note, we forced annotators to choose between struggle and non-struggle, instead of allowing a neutral option. This avoided centring strategies on ambiguous judgements and allowed easy binarisation of annotations during early pilot approaches that classed only struggle presence/absence. Throughout our research, we found that indicating the degree of struggle is necessary to address significant challenges in classifying fine-grained features.
As self-reporting levels of struggle can be unreliable and provide only a single judgement, we focus on a balance between expert and naive observers. We collected two distinct types of struggle determination annotations: naive ‘voter labels’ obtained from Amazon Mechanical Turk (MTurk) participants, and ‘expert annotations’ (or ‘golden annotation’ (GA)) obtained from the author who collected the video dataset, who repeatedly viewed them during analysis, and annotated action and error labels in the EPIC-Tent dataset. For each 10s video clip, there is only one expert annotation label, and there are multiple voters’ labels that form a frequency distribution. This distribution consists of 15 or 20 voter labels. To extract a single struggle label from this distribution, we take the mode. Since voter labels and expert annotation come from distinct groups of individuals with differing prior knowledge, label distributions can vary across all video segments for each dataset. Consistency between the voters’ modes and expert annotations indicates that the struggle-level annotations of the videos are more reliable. In addition, the standard deviation of the voters’ distributions provides a measure of the voter consistency or label ambiguity for a segment (smaller values indicate greater consistency, i.e. less label ambiguity). We explore these annotations in Section  3.2.
3.1.3 Expert Annotator Qualifications and Training
Unlike the MTurk workers who only viewed 10-second clips for each task, the expert, as a paper author, had previously reviewed the full videos (multiple times) during data annotation and quality control. This exposure provided them with a deeper understanding of the task context and its specific challenges. Additionally, the expert had prior experience coding actions, objects, and errors within the ‘tent’ dataset, further honing their observation and annotation skills relevant to the 10-second clips.
We acknowledge the inherent ambiguity in struggle determination, similar to skill assessment. Much like Olympic judges offering slightly different scores for a performance of exceptional skill, we anticipate slight disagreement in struggle determination. In our dataset of short clips, struggle typically manifests as motor errors, such as abrupt stops, repetitive attempts without success, object drops, or awkward manipulation. We believe that basic motor struggle is readily detectable by most human observers based on the fluidity of movement. However, the expert’s advantage lies in their extensive experience making such judgments, coupled with their unique knowledge of the full videos and specific task goals.
Figure 2: Frequency distributions of struggle labels for each split. The three graphs display the distribution of the four-way struggle labels for the activities Pipes-Struggle (left), Tent-Struggle (middle), and Tower-Struggle (right), respectively.
Figure 3: A summary of the number of video samples corresponding to participants for each split. The three graphs display participant IDs and the number of video segments for the activities Pipes-Struggle (left), Tent-Struggle (middle), and Tower-Struggle (right), respectively.
Figure 4: Comparison of the distributions. The first three graphs on the left display the frequency distribution of the struggle labels including expert annotation (EA), voters mode (Mode), and voters labels distributions (Voters), for the three activities. The three graphs on the right display the histogram with smoothed kernel density estimate (KDE) of the standard deviation (StdDev) of the voters’ labels for the three activities.
Table 2: Statistics for the three activities: the number of videos altogether and in each split, percentage of consistent videos in binary labels, percentage of consistent videos in four-class labels.
Tasks #Videos (Split 1/2/3/4) %Binary Consistent %Four-way Consistent
Pipes 1011 (275/229/244/263) 80.61% 45.40%
Tent 585 (138/147/146/154) 72.48% 42.91%
Tower 236 (65/54/67/50) 84.32% 52.12%
3.1.4 MTurk Annotator Validation
We recognize that struggle determination and its annotation is an ambiguous task, which is part of the challenge and research value for this problem. However, we still want to ensure annotation by crowd-sourcing is as valid as possible and that the votes and MTurkers can be trusted. We use the following assumptions in this ambiguous undertaking: 1) we assume the expert annotator is trustworthy, 2) ambiguity between sides of struggle (e.g. definitely struggling and slightly struggling and similarly for non-struggling) is acceptable and expected, 3) If a Mturker is untrustworthy in one of the assignments, all results by that MTurker is questionable and best to remove. Based on these assumptions we implemented the following validation checks:
  • If all votes for all videos assigned to the MTurker on an assignment are the same, all results from all assignments by that MTurker are rejected.
  • We also identified 6 validation videos using three definitely struggling and three definitely non-struggling videos as labelled by the expert annotator. These videos were further checked for unanimous agreement by simple inspection by the authors.
  • These 6 validation videos were included as part of the MTurk assignments and for every MTurker we calculate the level of disagreement w.r.t. the validation videos and the expert’s votes. If the MTurker’s votes are on the same side of struggle w.r.t the expert’s vote, i.e. (struggling (3,4) or non-struggling (1,2)) this does not count as a disagreement. The MTurker is rejected if the ratio of disagreement w.r.t. the validation videos is $>0.33$, i.e. three or more out of 6 disagreements for the validation videos.
3.2 Struggle Statistics
The three subsets of Our struggle determination dataset, Tent-Struggle, Pipes-Struggle, and Tower-Struggle, contain 585, 1011, and 236 annotated 10s videos, respectively, totalling  5.09 hours recorded with  725,100 frames. We further divide each of the subsets into four splits for four-fold cross-validation such that the video samples from the same participants are within the same split and the struggle label distributions are similar across the four splits (see Figure 2 for more details). The final split of participants and the corresponding number of video segments they contributed are shown in Figure 3. We note that some participants have contributed a significantly higher number of video segments than the other participants.
When considering struggle labels for video samples, there are three types of labels to take into account: a) Expert annotation—the struggling label given by the expert annotator; b) Voters’ mode—the statistical mode’s struggle value the voters give; and c) Voters’ label distribution—composed of the frequencies of voters’ labels on each of the struggle levels for each video sample.
Table 2 shows the statistics for the three subsets, including the total number of videos, the number of videos per split and the proportion of videos whose voters’ mode label is consistent with expert annotations.
Figure 4 shows both the overall frequencies for each struggle label and the histograms of standard deviations that indicate the label ambiguity for each dataset. The histograms of standard deviations tell us that the Tent-Struggle dataset has the most ambiguous voter labels with a standard deviation between 0.9 and 1. This is followed by the Pipes-Struggle dataset and then the Tower-Struggle dataset, which has the least ambiguous labels. These ambiguities reflect the degree of variation innate to each task with a well-defined task such as Towers of Hanoi having less ambiguity and assembling the non-rigid tent outdoors having the highest. The variation of ambiguity is a further element of interest for struggle-aware systems.
4 Experiments
Table 3: Results for the classification task. Classification accuracy for binary classification, four-way classification and the four-way to binary classification.
Dataset Pipes-Struggle Tent-Struggle Tower-Struggle
Top-1 Accuracy (%) Top-1 Accuracy (%) Top-1 Accuracy (%)
Models Binary Four-Way Four-to-Bin Binary Four-Way Four-to-Bin Binary Four-Way Four-to-Bin
Random 53.38 28.62 53.38 50.59 27.12 50.59 67.86 40.45 67.86
Majority Class 62.60 39.60 62.60 54.91 35.79 54.91 79.74 57.56 79.74
Voters’ Baseline 80.56 45.31 80.56 72.63 43.12 72.63 84.41 51.86 84.41
TSN [58] 74.54 48.59 71.43 68.05 38.23 61.24 82.59 59.49 78.68
SlowFast-R50 [16] 78.01 50.77 75.73 69.56 39.48 65.19 88.24 63.92 80.27
SlowFast-gMLP 78.59 49.70 74.49 68.00 39.52 64.68 87.86 66.40 83.17
VTN-SlowFast-ViT 77.80 50.83 72.48 68.53 40.11 60.86 88.11 64.03 82.54
We introduce three tasks for assessing struggle determination. First, struggle classification includes binary and four-way classifications using expert labels. In the four-way classification, we use the original struggle labels ranging from 1 to 4 corresponding to definitely non-struggle to definitely struggle. In binary classification, we binarize the label of struggling versus non-struggling by grouping 1&2 votes and 3&4 votes. Secondly, regression fits to the degree of struggle. Here we treat the struggle level labels from the experts as continuous real numbers ranging from 1 to 4. The third task is label distribution learning. Given the video samples, we train a deep model that can predict the frequency distribution over the four struggle levels while using the frequency distributions of the voters’ annotations for each video segment as the target. In the rest of this section, we present our baseline methods for end-to-end training, followed by the baseline experiment implementation details and results. We train and test our deep models using two Titan X (Pascal) 12GB GPUs or two GTX 1080Ti 11GB GPUs across all the experiments.
4.1 Baseline Methods
We implement three different deep architectures in our experiments: (1) 2D-ConvNet represented by the Temporal Segment Networks (TSN) [58], using the implementation from [63]. (2) 3D-ConvNet represented by the SlowFast Networks [16] based on ResNet50 (SlowFast-R50) [13], and (3) hybrid architectures of 3D-ConvNets and Vision Transformers [8], referred to as VTN-SlowFast-ViT and SlowFast-gMLP. The last two deep architectures implement the SlowFast Networks as the backbone to extract spatial-temporal features from the video sequence and are followed by different attention mechanisms such as a Vision Transformer layer [8] encoder block implemented based on layer [35] (VTN-SlowFast-ViT); or a gMLP layer [30] implemented from [34] (SlowFast-gMLP). Specifically, given a dataset which has a set of $K$ videos $P=\{p_{i},1\leq p_{i}\leq K\}$. We sample $N$ frames from each video sample $p_{i}$ and feed them into the SlowFast backbone $\Phi(\cdot)$, resulting in $N/8$ feature vectors that are then fed into the Vision Transformer encoder or a gMLP layer. This is followed by an MLP head for specific tasks, such as classification, regression, and label distribution learning. We use four-fold cross-validation to train and test these deep models for all tasks. We freeze the parameters in the backbone layers during training to overcome overfitting, which means linear probing for the first two deep architectures and only train the Vision Transformer-based layers for the hybrid architectures.
Table 4: Regression task results. Mean Squared Error (MSE) and Mean Absolute Error (MAE).
Activities Models Regression & Classification
[…]
Pipes-Struggle Voters’ Baseline 0.890 0.654 80.56% 45.31%
TSN [58] 0.798 0.725 72.42% 45.02%
SlowFast-R50 [16] 0.645 0.655 76.89% 45.28%
VTN-SlowFast-ViT 0.648 0.650 77.30% 48.56%
Tent-Struggle Voters’ Baseline 1.031 0.713 72.63% 43.12%
TSN [58] 0.888 0.768 68.27% 40.95%
SlowFast-R50 [16] 0.850 0.760 69.06% 41.63%
VTN-SlowFast-ViT 0.868 0.770 66.39% 40.69%
Tower-Struggle Voters’ Baseline 0.726 0.561 84.41% 51.86%
TSN [58] 0.935 0.748 85.35% 58.01%
SlowFast-R50 [16] 0.793 0.690 86.89% 56.00%
VTN-SlowFast-ViT 0.723 0.673 87.86% 61.73%
4.2 Classification Task
4.2.1 Implementation Details
We conducted two classification experiments: binary classification, in which we classify only struggle and non-struggle, and four-class classification, in which we predict the level of struggle. In either case, the model predicts the probability for each struggle level and is supervised by the Cross-Entropy (CE) loss. We use top-1 accuracy as an evaluation metric.
During training, we uniformly sample 128 frames from each video. The input frames are randomly cropped from $256\times 456$ to $256\times 256$. We also implement a horizontal flip at a probability of 0.5 during the training stage. In the test stage, we perform 5 random crops and ensemble the predictions. For TSN [58], we implement the BN-Inception base architecture with the average consensus for the output from each of the segments. We use ImageNet [45] pre-trained features and only train the fully-connected layer for classification. To overcome overfitting, we set the dropout rate to 0.5. We train for 65 epochs using mini-batch stochastic gradient descent with a momentum of 0.9 and weight decay 5e-4. The learning rate starts from 0.0001 and decreases 10×10$\times$10 × in epochs 30 and 60. For the SlowFast model and its variants we use the SlowFast-R50 and load the parameters pre-trained on Kinetics-400 [25]. During training, we freeze the parameters in the backbone and train the rest of the MLP layers. We use the same optimizer but the weight decay is set to 1e-5. We train these deep models for 70 epochs and implement a cyclic cosine decay learning rate scheduler with fixed cycle restart intervals [2] in which the learning rate starts from 0.001 and decreases 100 times in 50 epochs and followed by two restart intervals for 10 epochs each with the restart learning rate increase 10 times on the minimum value and decrease back to the minimum in every 10 epochs.
Table 5: Label distribution learning results. Mean Absolute Error (MAE) and Spearman’s Rank Correlation (Spearman’s Rho).
Activities Models Label Distribution Learning
[…]
Pipes-Struggle TSN [58] 0.13 0.5120 79.72% 46.01%
SlowFast-R50 [16] 0.12 0.5614 79.84% 48.79%
VTN-SlowFast-ViT 0.12 0.5689 80.30% 46.47%
Tent-Struggle TSN [58] 0.17 0.2307 54.04% 28.41%
SlowFast-R50 [16] 0.14 0.2887 57.07% 31.09%
VTN-SlowFast-ViT 0.13 0.3318 63.29% 34.95%
Tower-Struggle TSN [58] 0.24 0.4168 71.06% 37.06%
SlowFast-R50 [16] 0.15 0.4936 74.86% 40.44%
VTN-SlowFast-ViT 0.13 0.5689 77.58% 50.20%
4.2.2 Classification Results
The classification task results are shown in Table 3. The first three rows are naive baselines: (1) random performance using the frequency of classes as the probability; (2) majority class in percentage; and (3) voter’s baseline in which we use the mode of the voters’ to calculate the averaged accuracy compared with the expert annotations. In general, TSN has the lowest accuracy, because it doesn’t consider temporal relationships compared to other models based on 3D-CNNs. We can also see that there remains a gap between the voters’ baseline accuracy and the best performance of our deep models. We highlight the importance of temporal modelling in further ablations in Section 5. Finally, we note that the three 3D-ConvNets have similar accuracy, which means there is little benefit gained from implementing additional MLP/Transformer-based architectures for integrating temporal information in comparison with simply using global average pooling (GAP) over the temporal sampled segments. We hypothesise this is because we evaluate our models on short video sequences so the potential power of the MLP/Transformer layers is limited.
4.3 Regression Task
4.3.1 Implementation Details
In the regression task, the labels for struggle are treated as continuous scores indicating the degree of struggle, ranging from 1 (definitely non-struggle) to 4 (definitely struggle). We add a fully connected layer that takes the feature vectors as input and outputs a regression score. The predicted regression score is supervised by Mean Squared Error (MSE) together with target regression labels. The evaluation metrics include Mean Squared Error (MSE) and Mean Absolute Error (MAE) for regression and top-1 binary or four-way classification accuracy for classification. The hyperparameter settings of the optimization process are the same as those in the classification task.
4.3.2 Regression Results
Regression results are shown in Table 4. To obtain the classification results, we quantize the regression score into binary and four-way classification labels using evenly divided thresholds and then calculate classification accuracy. VTN-SlowFast-ViT generally has the lowest MSE and MAE and the highest classification accuracy rate on activities pipes and towers, followed by SlowFast-R50 and then TSN. We find that there is a greater variability of model performance when using continuous labels compared to the classification task. However, this trend does not hold on the Tent-Struggle dataset where the model VTN-SlowFast-ViT does not outperform the SlowFast-R50, either on regression metrics or on the converted classification metrics. This may be because the dataset tent is the most complex for struggle determination—it involves outdoor tent assembly scenes with five different sub-actions, so it is hard to optimize the extra ViT layer in the model VTN-SlowFast-ViT by only supervising a regression output.
4.4 Label Distribution Learning Task
4.4.1 Implementation Details
In label distribution learning, we train a deep model to predict voters’ struggle label frequency distribution given video samples. In our implementation, the final layer output of the model has four nodes corresponding to each of the four struggle levels ranging from 1 (definitely non-struggle) to 4 (definitely struggle) followed by a SoftMax layer so that each node’s output value ranges from 0 to 1 with the summation equaling to one. We optimize the Kullback–Leibler (KL) divergence loss function to align the predicted distribution with the frequency distribution of the voters’ struggle labels for each video sample. For evaluation, we calculate Spearmans’ Rank Correlation (Spearmans’ Rho) between the predicted frequency distribution and the voters’ struggle label frequency distribution so that the value ranges from -1 to 1 with a higher score indicating that the predicted distribution better aligns with its target. We also calculate the Mean Absolute Error (MAE) between the predicted distribution and the target distribution. For converted classification, we calculate the accuracy rate between the predicted highest frequency label and the voters’ mode label. The hyper-parameter settings are the same as those in the previous tasks.
Figure 5: Importance of Multiple Frames. The three bar charts on the left depict a binary classification scenario for the Pipes-Struggle, Tent-Struggle, and Tower-Struggle activities, while the three bar charts on the right depict a four-way classification scenario in the same order. Each group of bars represents the base model (blue); evaluating using a single frame (orange); and training and evaluating using a single frame (green). Error bars show the standard deviation (StdDev) across the sampled frames from three proportions (25%, 50%, and 75%) of the video frames counted starting at the beginning.
Figure 6: Importance of Temporal Ordering. The three bar charts on the left depict a binary classification scenario for the Pipes-Struggle, Tent-Struggle, and Tower-Struggle activities, while the three bar charts on the right depict a four-way classification scenario in the same order. Each group of bars represents the input frames in the correct order (blue); the evaluating using shuffled frames (orange); and the training and evaluating using shuffled frames (green).
4.4.2 Distribution Results
The results for label distribution learning are shown in Table 5. Due to the continuous nature of MAE and Spearman’s Rho evaluation metrics, it is possible to demonstrate discrepancies in the evaluation metrics between the model TSN, SlowFast-R50, and VTN-SlowFast-ViT that lead to the same trend of discrepancies in the corresponding converted accuracy rate. On all evaluation metrics we implemented in this task, the 3D-ConvNet SlowFast outperforms the 2D-ConvNet TSN, and the VTN-SlowFast-ViT is typically superior to the SlowFast. The spatial-temporal features of SlowFast appear to be more useful than the 2D spatial features in TSN, and the additional ViT layer provides a better fit for the mode label distributions when compared to vanilla SlowFast-R50. The classification results shown in Table 5 are obtained from the predicted voters’ label frequency distribution. we convert the frequency distribution into class labels by selecting the label with top-1 frequency as a four-way classification label and combining the predicted frequency of 1 with 2, and 3 with 4 to get binary classification frequency for binary classification. Therefore, the converted classification results can also be considered a classification baseline based on voters’ mode labels.
5 Ablation Study
In this section, we investigate the importance of temporal modelling and motion understanding for the struggle determination task. Specifically, we conduct two experiments: Firstly, how much information can be gleaned from a single frame, and secondly, how important is the temporal ordering of frames.
5.1 Multiple Frames
To show the necessity of using multiple frames, we conducted experiments using one frame and made a comparison with the original multi-frame setting. We first loaded the model parameters trained using 128 frames which is the standard setting for our experiments and then ran only the test stage by loading one frame from each of the video sequences. The one frame is selected from proportions of 25%, 50%, and 75% respectively from the beginning of each video segment. Note that we duplicated the sampled one frame to meet the minimum number of input frames required by 3D-ConvNets, which is 8 for SlowFast, 128 for SlowFast-gMLP and VTN-SlowFast-ViT. To maintain consistency between the training and testing stages, we conducted additional experiments by randomly selecting one frame from each video during the training stage and one frame from the same proportions during the testing stage. The results for the two cases described above are shown in Figure 5. We report the one-frame test accuracy using average accuracy from proportions of 25%, 50%, and 75% respectively. There is a significant decrease in accuracy across all of the models we tested on all three activities. And because the video samples in the dataset are randomly trimmed into 10-second segments, the accuracy rates for the frame sampled from 25%, 50%, and 75% are expected to be similar, which means a small standard deviation (StdDev) of accuracy close to 0 in the bar plots. Even though we trained the model using a single frame, the test results still showed a notable decrease in accuracy, despite a slight increase compared to testing the models with only one frame. These results indicate that no sufficient visual information is provided for the deep learning models to rely on just one frame to determine whether a person is struggling. Therefore, effectively capturing the useful temporal motion features is crucial for our struggle determination task. A sequence of frames can typically reveal anomalous hand motions such as ‘unsmooth’ motions, stopping, and repeating that indicate the person did not have things under control. A single frame cannot represent these features. To test this idea more rigorously, however, it is important to present the models with a sequence of frames that do not contain useful temporal motion features.
Figure 7: Visualization results of the activation heatmaps and temporal attention scores. Visualization results for both non-struggle (upper box) and struggle (bottom box) are shown from activities: Pipes-Struggle (top), Tent-Struggle (middle), and Tower-Struggle (bottom) respectively with each row corresponding to the frames sampled from a video sample that is fed into the slow pathway and down-sampling into 8 frames. The heatmaps are the Grad-CAM++ [5] visualization and the green bars below are the colourmap visualizations for the temporal attention scores of the corresponding video segments. The intensity of the green indicates the value of the attention score range from 0 to 1. The closer to 1, the darker the green colour.
5.2 Temporal Ordering
We ran a shuffle-frame test to show the importance of temporal ordering. We did not consider TSN because it only considers 2D spatial features and calculates the average of the output logits of all the input frames, making it insensitive to the order of the frames. We first loaded the model parameters pre-trained using 128 frames in the right order. Then we loaded the same number of 128 frames from each of the video samples and shuffled them randomly so that we disrupted the correct order of the video frames. We ran the model in inference mode ten times and calculated the average accuracy. As shown in Figure 6, the accuracy rate decreased significantly to the random level, which means the models are very sensitive to the frame orders. To make a fair comparison, we conducted another series of experiments in which we also trained the models using shuffled frames and then reported the test shuffle-frame accuracy. Figure 6 shows that after training using shuffled frames, the models can perform much better than when solely conducting the test stage, but still lower than the results when using the frames in the correct order. Our illustration suggests that the deep models learned key temporal dependencies of video sequences when they were trained using the correct order of frame sequences; as a result, when we test the models using shuffled frames, their accuracy approaches random performance.
6 Attention Visualization
In this section, we provide visualization towards what the models are focusing on via their attention weights/activation maps [5]. This is key to ensuring that the models are learning how to detect the struggles of participants. Specifically, we visualize the activation maps from the backbone and the attention scores in the self-attention layer of the VTN-SlowFast-ViT model. In Figure 7 we show the visualization heatmaps from the final convolutional layer in the backbone network using Grad-CAM++ [5] and the attention score visualized using green bars. We present one sample from each of the three activities: Pipes-Struggle, Tent-Struggle, and Tower-Struggle. We can observe that the model chooses to focus on the hands and active objects in the frames, despite occasionally being noisy and focusing on the background. The corresponding green bars below show that the model is giving each frame varying amounts of attention. For example, in the Pipes-Struggle visualised frames, the heatmaps highlight the hand of the person holding the pipes as well as the pipes that are lying in the background. Similar to this, in the Tent-Struggle examples, the model focuses on the tent’s poles and the hands of the person attempting to put it up, as shown by the heatmaps. In the Tower-Struggle examples, the heatmaps mainly lie on the hands and the Hanoi puzzle blocks. Due to the limit of the 10-second video clips that cannot feature the entire action, the green bars are typically not aligned with explicit actions but still tend to indicate a preference for specific frames within the sampled video sequence.
7 Discussion and Conclusions
In this paper, we have shown and discussed how struggle determination is an important component of human video understanding and a key element for building assistive systems. We have introduced a dataset including three activities for the struggle determination task, with a total of 5.09 hours of video.
Our results reveal opportunities for improvement in the four-way classification task aimed at finer-grained assessment of struggle. The similar accuracy rates between four-way to binary conversion and binary classification indicate that the main challenge lies in distinguishing the degree of struggle. This is key for useful and dependable assistive systems. The tent assembly task has the lowest accuracy ($40\%$), for the four-way task, supporting our hypothesis that the ambiguity on task judgement for the tent pitching activity is harder than for the other activities that have more structure. The regression and label distribution learning tasks have encouraging results, demonstrating their efficacy in distinguishing model differences and determining struggle in different ways. In particular, the VTN-SlowFast-ViT model overall performs well in the label distribution learning tasks, yet albeit better than random, it still achieves low performance ($34\%$) for the four-way Tent-Struggle.
Several limitations and challenges remain: (1) Limited background scene variations across the dataset, consisting mostly of two indoor scenes and one outdoor scene for tent assembly. (2) We have only video-level ‘weak’ labels for struggle, finer temporal localisation and further task knowledge could help the model to improve performance. (3) Current pre-trained model parameters based on coarse-grained video action recognition may be sub-optimal for fine-grained struggle determination. (4) By using randomly trimmed 10-second video segments, we may overlook certain types of struggling that are more obvious in the entire video and can only focus on ‘low-level’ struggling detection. (5) A better understanding and validation of what is an appropriate human (or statistical) baseline is likely needed to validate model performance.
In conclusion, this paper aims to motivate and support further work on determining the level of struggle in video. A new annotated dataset for struggling with three real daily living activities is introduced. This data is with video-level annotations from both crowd voters and human experts. Three research challenges, namely classification, regression, and label distribution learning, are presented along with definitions and baseline results using various deep architectures. We hope this research provides valuable insights for additional work on fine-grained egocentric video understanding and supports new methods and directions to benchmark assistive systems.
[…]
Dataset Access Our struggle determination dataset can be found via this URL:
https://github.com/FELIXFENG2019/Struggle-Determination. (Temporary during review, archival repository will be provided at acceptance.)
[…]
Acknowledgments We are grateful to Prof. Dima Damen’s contributions for early discussions and involvement in the initial investigation of this work. Shijia Feng is supported by a scholarship from the China Scholarship Council (No.202109210007). Initial research, data capture and annotation supported by EPSRC grant GLANCE (EP/N013964/1).
Appendix A Related Datasets
Table 6: Comparison of related datasets. Note that AR is short for Action Recognition. AQA: Action Quality Assessment. VAD: Video Anomaly Detection.
Dataset Year Source Viewpoint Tasks Available Annotations Size
Coarse-grained AR
HMDB51 [26] 2011 Movies and online Third-person Human motion recognition 51 action categories 6,766 video clips
UCF101 [47] 2012 YouTube Third-person Action recognition 101 action classes Over 13k clips
ActivityNet [12] 2015 YouTube Third-person Human activity understanding 200 activities 19,994 untrimmed videos
Kinetics [25] 2017-2020 YouTube Third-person Action recognition 400/600/700 human action classes Up to 650,000 video clips
Fine-grained AR
SSV2 [19] 2017 In-person recording First-person Fine-grained understanding of human-object interaction 174 labels for different types of hand-object interactions 108,499 videos
Diving48 [27] 2018 Web videos Third-person Fine-grained diving action recognition 48 dive classes About 18k videos
FineGym [46] 2020 YouTube Third-person Fine-grained action recognition 15 set classes, 530 element classes, 4,883 instances, 32,697 sub-instances 303 competition records
AQA
Olympic Scoring [42] 2017 Existing datasets and online Third-person Olympics sports score assessment score ranges from 0 or 20 to 100 696 videos
JIGSAW [18] 2017 JHU and Sunnyvale, CA. ISI First-person Surgery skill assessment 15 gestures and 5-point Likert scale rating score 206 videos
AQA-7 [41] 2019 YouTube Third-person Action quality assessment across seven sports AQA scores range from 6.72 to 104.88 altogether across seven sports 1189 videos
MTL-AQA [43] 2019 Online Third-person Multi-task learning including fine-grained action recognition, commentary generation, AQA score estimation 16 classes of events, 5,748 commentaries, and AQA scores range from 0 to 100 1,412 video samples
FineDiving [62] 2022 YouTube Third-person Coarse- to fine-grained diving action procedures and AQA 52 action types and 29 sub-action types with temporal boundaries, along with judges’ scores, and 23 difficulty degree types. 3,000 video samples
Skill Determination
EPIC-Skill2018 [9] 2018 Existing dataset and in-person recording First-person Pairwise skill ranking including four tasks Lists of video 1, video 2, and the better video in the pair. 196 video samples
BEST [10] 2019 In-person recording First-person Pairwise skill ranking including five tasks Lists of video 1, video 2, and the better video in the pair. 500 videos
TikTok Dance [21] 2023 Participants on TikTok Third-person Pairwise dance skill ranking tasks including 12 dance challenges Video pairs with the winning (better) video. 240 videos
VAD
UCSD-Peds [36] 2013 Surveillance camera Third-person Determine videos with abnormal actions Binary per-frame flag indicator and pixel-level binary masks for abnormal regions 70 videos in Ped 1 and 28 videos in Ped 2
ShanghaiTech [31] 2018 Surveillance camera Third-person Predict future frames and determine abnormal events Frame-level masks and pixel-level masks for abnormal events 437 videos
UCF-Crime [50] 2018 Surveillance videos Third-person Determine videos with abnormal actions Video-level weak labels for anomalies 1,900 untrimmed videos
Struggle Determination
Ours 2023 Existing dataset and in-person recording First-person Determine struggle and degrees of struggle under three sub-tasks: struggle classification, struggle level regression, and struggle label distribution learning Video-level binary labels for struggle/non-struggle or four-scale struggle levels from 1 to 4. 1,832 trimmed video segments.
Table 6 contains a comparison of datasets in video understanding previously used for coarse-grained Action Recognition (AR), fine-grained AR, Action Quality Assessment (AQA), Skill Determination, Video Anomaly Detection (VAD) and Struggle Determination. These datasets are relevant to our struggle determination task as discussed in Section 2.
Action recognition datasets encompass both coarse-grained and fine-grained human actions. Deep models trained on large-scale action recognition datasets can be transferred to various downstream tasks due to the diverse scenarios and learned features. However, these datasets do not focus on scenes depicting struggling, particularly in the case of coarse-grained AR, which mainly includes common human actions and lacks scenarios for skills.
Action quality assessment datasets primarily annotate action quality scores for videos showcasing sports or surgical skills. However, lower action quality scores do not necessarily indicate struggling, and well-performed actions can still receive relatively low scores.
Skill determination datasets are more relevant to our struggle determination task as they often showcase different skills at different levels with hand-object interactions. However, the annotations in these datasets are in the form of pairwise rankings, indicating which skill is better than the other. Lower-ranked videos in these pairs may not guarantee that a person is struggling.
Video anomaly detection datasets mainly consist of surveillance videos capturing unusual events, such as fights or unintentional falls. These datasets are typically used for binary classification tasks to distinguish between anomaly and normal videos. While struggle determination also involves binary classification, it falls within the context of determining skills. Moreover, our goal is not only to differentiate struggle from non-struggle but also to determine the degree of struggle, which is crucial for developing a wearable intelligent system to provide appropriate assistance to the user.
In comparison, our struggle determination datasets encompass three tasks that demonstrate different types of skills in assembling pipes, pitching tents, and playing the Tower of Hanoi game. These datasets cover indoor and outdoor scenes and include both rigid and non-rigid objects. Additionally, we specifically use a struggle scoring scale, and thus, the annotations in our dataset are not limited to binary labels but are further fine-grained to indicate the degree of struggle. Aside from the focus on struggle, this distinguishes them from the class labels in AR, action quality scores in AQA, ranking labels in skill determination, and even the binary contrastive labels in VAD. Hence, it is essential to create a dataset with a focus on struggle annotations to train deep models for determining struggle and quantifying its degree.
Appendix B Dataset Collection Details
B.1 Instructions for Participants
During our data collection process, instructions were given to our in-person participants both verbally and written on paper. For the plumbing pipes task, there are two instruction diagrams illustrating how to assemble the pipes correctly, one version is easy to follow, shown in Figure 8, and another more difficult version, shown in Figure 9 so that the participants using a hard version of instructions are more likely to show signs of struggle. For the tent-pitching task, we used the paper instructions from the manufacturer as shown in Figure 10. For the Tower of Hanoi game task, the researchers only provided the instructions verbally introducing the basic rules of the game, such as only one disk can be moved at a time and a larger disk cannot be placed on top of a smaller one.
B.2 Additional Dataset Statistics
We add Figure 11 to display the struggle labels frequency distributions in the binary case (non-struggle vs struggle).
Appendix C Additional Experiment Results
C.1 Classification Task
We show more details of our struggle classification task about the top-1 accuracy rate for each of the four splits together with the mean and standard deviation (StdDev) over the accuracy of the four splits. Table 7, 8, and 9 show the additional results for binary classification, four-way classification, and four-way to binary classification, respectively.
We first give the full definitions of the three naive baselines: random, majority class, and voters’ baseline. Given a dataset $D$ (Pipes-Struggle, Tent-Struggle, or Tower-Struggle), $N$ is the number of videos. Let $V_{i}$ denote the $i$th video. The expert annotation for video $V_{i}$ is $y^{e}_{i}$ and the frequency distribution of the struggle labels given by the 15 or 20 voters $Dist_{i}$ is denoted as $Dist_{i}=\{f_{1},f_{2},f_{3},f_{4}\}$ where $f_{1}$ is the frequency of the voters label for struggle level 1 (definitely non-struggle), noted as $Dist_{i}(f_{1})$; $f_{2}$ is the frequency of the voters label for struggle level 2 (slightly non-struggle), noted as $Dist_{i}(f_{2})$; $f_{3}$ is the frequency of the voters label for struggle level 3 (slightly struggle), noted as $Dist_{i}(f_{3})$; $f_{4}$ is the frequency of the voters label for struggle level 4 (definitely struggle), noted as $Dist_{i}(f_{4})$, while $f_{1}+f_{2}+f_{3}+f_{4}=1$. Let the mode of the voters’ label for video $V_{i}$ be $y^{l}_{i}$. $y^{l}_{i}$ is the struggle label index (from 1 to 4 ) of $Mode(Dist_{i})$. Note that in the case of binary classification, we combine 1 and 2 to 0 (non-struggle), and 3 and 4 to 1 (struggle) as discussed in the main paper. Moreover, the label distributions over a certain split are calculated by averaging the expert annotation $y^{e}_{i}$ or the voters’ label distributions $Dist_{i}=\{f_{1},f_{2},f_{3},f_{4}\}$ across all the videos $V_{i}$ within the split, denoted as $Dist(y_{0})$ and $Dist(y_{1})$ respectively.
In this way, for the random baseline, we assume the random model only has the ability to predict each of the struggle levels in a probability that corresponds to the proportions of the expert annotations in the test split $Dist(y_{0})$, so that we can calculate the top-1 accuracy rate each time we run the random model. When the number of runs tends to be a large number close to infinity, the averaged top-1 accuracy rate over the number of runs is approaching a set percentage which is the random baseline accuracy. The majority class baseline is the percentage of the maximum entity in $Dist(y_{0})$ for a split. Finally, we illustrate the voter’s baseline. Given a video $V_{i}$, we define the Kronecker delta function:
$\delta_{y_{i}^{e}y_{i}^{l}}=\begin{cases}1&\text{if }y_{i}^{e}=y_{i}^{l}\\ 0&\text{if }y_{i}^{e}\neq y_{i}^{l}\end{cases}$ (1)
where $y_{i}^{e}$ is the expert annotation and $y_{i}^{l}$ is the mode of the voters’ labels. So the voters’ baseline accuracy is calculated as:
$\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\delta_{y_{i}^{e}y_{i}^{l}}\times 100\%$ (2)
where the $N_{j}$ is the number of videos in a split $j$.
C.2 Regression Task
We provide more experimental results on our struggle level regression task with the evaluation metrics of Mean Squared Error (MSE), Mean Absolute Error (MAE), and the top-1 accuracy rate of binary and four-way classification converted from the predicted struggle level regression scores. These evaluation metrics of various deep models on the three activities for split 1 to 4 are shown in Tables 10, 11, 12, and 13, respectively. The mean and standard deviation (StdDev) for these evaluation metrics for the four splits are shown in Table 14.
C.3 Label Distribution Learning Task
We provide more experimental results about voters’ struggle label distribution learning task with the evaluation metrics of Mean Absolute Error (MAE), Spearman’s Rank Correlation (Spearman’s Rho), and the top-1 accuracy rate of binary and four-way classification by selecting the struggle level classes with the highest frequency in the label distributions and using the voters’ mode label as ground truth. These evaluation metrics of various deep models on the three activities for split 1 to 4 are shown in Tables 15, 16, 17, and 18, respectively. The mean and standard deviation (StdDev) for these evaluation metrics for the four splits are shown in Table 19.
C.4 Further Ablation Details
We provide more experiment results of the ablation study we conducted using one frame as input and shuffled frame orders, see Table 20 and Table 21.
Appendix D Figures and Tables
Figure 8: The diagram of the easy double sink layout.
Figure 9: The diagram of the difficult double sink layout.
Figure 10: An OCR photocopy of the assembly instructions for the tent we use.
Figure 11: Comparison of the distributions for binary struggle labels. The three graphs display the frequency distribution of the binary struggle labels (0 for non-struggle, 1 for struggle) including expert annotation (EA), voters mode (Mode), and voters labels distributions (Voters), for the three activities: Pipes-Struggle, Tent-Struggle, and Tower-Struggle, respectively.
Table 7: Experiment results for binary classification task over all splits.
Activities Models Top1 Accuracy Rate (%) - Binary Classification
Split 1 Split 2 Split 3 Split 4 Mean±StdDev
EPIC-Pipes Random 55.00% 51.60% 51.93% 54.98% 53.38±1.87%
Majority Class 65.82% 58.95% 59.84% 65.78% 62.60±3.72%
Voters’ Baseline 83.27% 79.91% 80.33% 78.71% 80.56±1.68%
TSN [58] 75.27% 76.86% 74.18% 71.86% 74.54±2.10%
SlowFast-R50 [16] 77.09% 75.11% 80.74% 79.09% 78.01±2.44%
SlowFast-gMLP 78.91% 76.42% 80.33% 78.71% 78.59±1.62%
VTN-SlowFast-ViT 77.82% 75.55% 79.51% 78.33% 77.80±1.66%
EPIC-Tent Random 50.37% 51.02% 50.94% 50.03% 50.59±0.47%
Majority Class 54.35% 57.14% 56.85% 51.30% 54.91±2.71%
Voters’ Baseline 79.71% 73.47% 68.49% 68.83% 72.63±4.54%
TSN [58] 74.64% 70.07% 66.44% 61.04% 68.05±5.75%
SlowFast-R50 [16] 76.81% 72.11% 64.38% 64.94% 69.56±5.98%
SlowFast-gMLP 73.19% 73.47% 63.01% 62.34% 68.00±6.16%
VTN-SlowFast-ViT 73.91% 72.79% 65.07% 62.34% 68.53±5.70%
EPIC-Tower Random 71.88% 65.43% 70.59% 63.53% 67.86±4.01%
Majority Class 83.08% 77.78% 82.09% 76.00% 79.74±3.39%
Voters’ Baseline 81.54% 87.04% 85.07% 84.00% 84.41±1.98%
TSN [58] 86.15% 79.63% 86.57% 78.00% 82.59±4.41%
SlowFast-R50 [16] 89.23% 85.18% 92.54% 86.00% 88.24±3.36%
SlowFast-gMLP 89.23% 85.18% 91.04% 86.00% 87.86±2.75%
VTN-SlowFast-ViT 89.23% 85.18% 94.03% 84.00% 88.11±4.54%
Table 8: Experiment results for four-way classification task over all splits.
Activities Models Top1 Accuracy Rate (%) - Four-way Classification
Split 1 Split 2 Split 3 Split 4 Mean±StdDev
EPIC-Pipes Random 28.82% 29.65% 27.88% 28.12% 28.62±0.69%
Majority Class 40.36% 41.92% 38.11% 38.02% 39.60±1.63%
Voters’ Baseline 47.27% 46.29% 39.75% 47.91% 45.31±3.26%
TSN [58] 48.00% 53.71% 45.49% 47.15% 48.59±3.09%
SlowFast-R50 [16] 53.46% 50.66% 48.77% 50.19% 50.77±1.70%
SlowFast-gMLP 53.46% 48.04% 46.72% 50.57% 49.70±2.58%
VTN-SlowFast-ViT 50.54% 51.53% 48.77% 52.47% 50.83±1.37%
EPIC-Tent Random 26.71% 27.15% 28.87% 25.73% 27.12±1.14%
Majority Class 35.51% 36.05% 40.41% 31.17% 35.79±3.27%
Voters’ Baseline 50.72% 38.10% 47.95% 35.71% 43.12±6.35%
TSN [58] 40.58% 39.46% 40.41% 32.47% 38.23±3.35%
SlowFast-R50 [16] 45.65% 38.10% 39.73% 34.42% 39.48±4.05%
SlowFast-gMLP 41.30% 40.82% 36.99% 38.96% 39.52±1.70%
VTN-SlowFast-ViT 42.75% 40.14% 42.47% 35.06% 40.11±3.08%
EPIC-Tower Random 42.77% 38.54% 44.71% 35.76% 40.45±3.51%
Majority Class 60.00% 55.56% 62.69% 52.00% 57.56±4.10%
Voters’ Baseline 56.92% 59.26% 49.25% 42.00% 51.86±6.79%
TSN [58] 63.08% 53.70% 67.16% 54.00% 59.49±5.82%
SlowFast-R50 [16] 66.15% 64.82% 62.69% 62.00% 63.92±1.66%
SlowFast-gMLP 64.62% 64.82% 70.15% 66.00% 66.40±2.23%
VTN-SlowFast-ViT 67.69% 59.26% 67.16% 62.00% 64.03±3.54%
Table 9: Experiment results of four-way classification convert into binary classification over all splits.
Activities Models Top1 Accuracy Rate (%) - Four-way to Binary
Split 1 Split 2 Split 3 Split 4 Mean±StdDev
EPIC-Pipes Random 55.00% 51.60% 51.93% 54.98% 53.38±1.62%
Majority Class 65.82% 58.95% 59.84% 65.78% 62.60±3.22%
Voters’ Baseline 83.27% 79.91% 80.33% 78.71% 80.56±1.68%
TSN [58] 73.82% 71.52% 69.67% 70.72% 71.43±1.53%
SlowFast-R50 [16] 78.18% 73.36% 74.18% 77.19% 75.73±2.01%
SlowFast-gMLP 78.18% 72.93% 74.59% 72.24% 74.49±2.30%
VTN-SlowFast-ViT 72.73% 71.18% 73.77% 72.24% 72.48±0.93%
EPIC-Tent Random 50.37% 51.02% 50.94% 50.03% 50.59±0.41%
Majority Class 54.35% 57.14% 56.85% 51.30% 54.91±2.35%
Voters’ Baseline 79.71% 73.47% 68.49% 68.83% 72.63±4.54%
TSN [58] 67.39% 63.95% 62.33% 51.30% 61.24±6.02%
SlowFast-R50 [16] 72.46% 70.75% 63.01% 54.55% 65.19±7.10%
SlowFast-gMLP 67.39% 72.11% 57.53% 61.69% 64.68±5.54%
VTN-SlowFast-ViT 68.84% 63.27% 55.48% 55.84% 60.86±5.56%
EPIC-Tower Random 71.88% 65.43% 70.59% 63.53% 67.86±3.47%
Majority Class 83.08% 77.78% 82.09% 76.00% 79.74±2.94%
Voters’ Baseline 81.54% 87.04% 85.07% 84.00% 84.41±1.98%
TSN [58] 83.08% 74.07% 83.58% 74.00% 78.68±4.65%
SlowFast-R50 [16] 84.62% 70.37% 82.09% 84.00% 80.27±5.79%
SlowFast-gMLP 87.69% 75.93% 85.07% 84.00% 83.17±4.39%
VTN-SlowFast-ViT 86.15% 75.93% 88.06% 80.00% 82.54±4.84%
Table 10: Experiment results for regression task—split 1.
Activities Models Regression & Classification - Split 1
[…]
EPIC-Pipes TSN [58] 0.83 0.72 72.36% 45.82%
SlowFast-R50 [16] 0.64 0.67 77.82% 43.27%
VTN-SlowFast-ViT 0.68 0.66 77.45% 48.36%
EPIC-Tent TSN [58] 0.82 0.74 69.57% 42.03%
SlowFast-R50 [16] 0.75 0.72 73.91% 42.03%
VTN-SlowFast-ViT 0.78 0.74 68.84% 41.30%
EPIC-Tower TSN [58] 0.75 0.68 86.15% 58.46%
SlowFast-R50 [16] 0.77 0.70 87.69% 58.46%
VTN-SlowFast-ViT 0.70 0.67 87.69% 60.00%
Table 11: Experiment results for regression task—split 2.
Activities Models Regression & Classification - Split 2
[…]
EPIC-Pipes TSN [58] 0.64 0.65 76.42% 50.22%
SlowFast-R50 [16] 0.60 0.62 73.46% 48.47%
VTN-SlowFast-ViT 0.67 0.65 74.24% 49.78%
EPIC-Tent TSN [58] 0.94 0.78 73.47% 42.86%
SlowFast-R50 [16] 0.79 0.72 72.79% 45.58%
VTN-SlowFast-ViT 0.85 0.76 68.71% 39.46%
EPIC-Tower TSN [58] 0.93 0.76 85.19% 57.41%
SlowFast-R50 [16] 0.78 0.69 83.33% 51.85%
VTN-SlowFast-ViT 0.88 0.72 85.19% 59.26%
Table 12: Experiment results for regression task—split 3.
Activities Models Regression & Classification - Split 3
[…]
EPIC-Pipes TSN [58] 0.83 0.74 71.31% 42.21%
SlowFast-R50 [16] 0.61 0.63 78.69% 47.95%
VTN-SlowFast-ViT 0.56 0.61 80.33% 51.23%
EPIC-Tent TSN [58] 0.87 0.76 65.75% 43.84%
SlowFast-R50 [16] 0.82 0.75 68.49% 43.84%
VTN-SlowFast-ViT 0.87 0.77 63.70% 41.10%
EPIC-Tower TSN [58] 0.87 0.71 88.06% 64.18%
SlowFast-R50 [16] 0.49 0.54 92.54% 65.67%
VTN-SlowFast-ViT 0.58 0.62 92.54% 71.64%
Table 13: Experiment results for regression task—split 4.
Activities Models Regression & Classification - Split 4
[…]
EPIC-Pipes TSN [58] 0.89 0.79 69.58% 41.83%
SlowFast-R50 [16] 0.73 0.70 77.57% 41.44%
VTN-SlowFast-ViT 0.68 0.68 77.19% 44.87%
EPIC-Tent TSN [58] 0.92 0.79 64.29% 35.06%
SlowFast-R50 [16] 1.04 0.85 61.04% 35.06%
VTN-SlowFast-ViT 0.97 0.81 64.29% 40.91%
EPIC-Tower TSN [58] 1.19 0.84 82.00% 52.00%
SlowFast-R50 [16] 1.13 0.83 84.00% 48.00%
VTN-SlowFast-ViT 0.73 0.68 86.00% 56.00%
Table 14: Experiment results for regression task—results across all four splits including mean values and standard deviation (StdDev).
Activities Models Regression & Classification - Statistics (Mean±StdDev)
[…]
EPIC-Pipes TSN [58] 0.798±0.094 0.725±0.050 72.42±2.52% 45.02±3.38%
SlowFast-R50 [16] 0.645±0.051 0.655±0.032 76.89±2.02% 45.28±3.00%
VTN-SlowFast-ViT 0.648±0.051 0.650±0.025 77.30±2.16% 48.56±2.36%
EPIC-Tent TSN [58] 0.888±0.047 0.768±0.019 68.27±3.57% 40.95±3.46%
SlowFast-R50 [16] 0.850±0.112 0.760±0.053 69.06±5.05% 41.63±3.99%
VTN-SlowFast-ViT 0.868±0.068 0.770±0.025 66.39±2.40% 40.69±0.72%
EPIC-Tower TSN [58] 0.935±0.161 0.748±0.061 85.35±2.19% 58.01±4.32%
SlowFast-R50 [16] 0.793±0.227 0.690±0.103 86.89±3.66% 56.00±6.72%
VTN-SlowFast-ViT 0.723±0.107 0.673±0.036 87.86±2.85% 61.73±5.92%
Table 15: Label distribution learning experiment results on split 1.
Activities Models Label Distribution Learning - Split 1
[…]
EPIC-Pipes TSN [58] 0.12 0.4922 80.00% 39.27%
SlowFast-R50 [16] 0.11 0.5484 80.00% 44.73%
VTN-SlowFast-ViT 0.11 0.5316 78.55% 38.55%
EPIC-Tent TSN [58] 0.16 0.2743 52.17% 29.71%
SlowFast-R50 [16] 0.13 0.3658 71.74% 37.68%
VTN-SlowFast-ViT 0.13 0.3785 64.49% 36.96%
EPIC-Tower TSN [58] 0.30 0.4881 73.85% 35.38%
SlowFast-R50 [16] 0.14 0.4938 73.85% 30.77%
VTN-SlowFast-ViT 0.13 0.5709 76.92% 44.62%
Table 16: Label distribution learning experiment results on split 2.
Activities Models Label Distribution Learning - Split 2
[…]
EPIC-Pipes TSN [58] 0.13 0.4771 78.17% 41.92%
SlowFast-R50 [16] 0.12 0.5012 77.73% 45.41%
VTN-SlowFast-ViT 0.12 0.5113 78.17% 42.79%
EPIC-Tent TSN [58] 0.18 0.1983 50.34% 21.09%
SlowFast-R50 [16] 0.14 0.2048 53.74% 27.21%
VTN-SlowFast-ViT 0.13 0.2941 63.95% 38.10%
EPIC-Tower TSN [58] 0.17 0.4137 72.22% 37.04%
SlowFast-R50 [16] 0.14 0.5532 75.93% 48.15%
VTN-SlowFast-ViT 0.12 0.5879 77.78% 57.41%
Table 17: Label distribution learning experiment results on split 3.
Activities Models Label Distribution Learning - Split 3
[…]
EPIC-Pipes TSN [58] 0.12 0.5605 82.38% 55.33%
SlowFast-R50 [16] 0.13 0.5940 79.51% 53.69%
VTN-SlowFast-ViT 0.12 0.6202 81.97% 52.05%
EPIC-Tent TSN [58] 0.17 0.2733 63.01% 34.93%
SlowFast-R50 [16] 0.13 0.3282 54.11% 32.19%
VTN-SlowFast-ViT 0.13 0.3536 63.01% 33.56%
EPIC-Tower TSN [58] 0.24 0.4273 70.15% 35.82%
SlowFast-R50 [16] 0.15 0.4739 71.64% 32.84%
VTN-SlowFast-ViT 0.13 0.4985 77.61% 44.78%
Table 18: Label distribution learning experiment results on split 4.
Activities Models Label Distribution Learning - Split 4
[…]
EPIC-Pipes TSN [58] 0.13 0.5181 78.33% 47.53%
SlowFast-R50 [16] 0.12 0.6020 82.13% 51.33%
VTN-SlowFast-ViT 0.12 0.6125 82.51% 52.47%
EPIC-Tent TSN [58] 0.17 0.1769 50.65% 27.92%
SlowFast-R50 [16] 0.14 0.2560 48.70% 27.27%
VTN-SlowFast-ViT 0.14 0.3011 61.69% 31.17%
EPIC-Tower TSN [58] 0.24 0.3382 68.00% 40.00%
SlowFast-R50 [16] 0.15 0.4534 78.00% 50.00%
VTN-SlowFast-ViT 0.12 0.6183 78.00% 54.00%
Table 19: Label distribution learning experiment results—results across all the four splits including mean values and standard deviation (StdDev).
Activities Models Label Distribution Learning - Statistics (Mean±StdDev)
[…]
EPIC-Pipes TSN [58] 0.13±0.01 0.5120±0.0365 79.72±1.96% 46.01±7.10%
SlowFast-R50 [16] 0.12±0.01 0.5614±0.0466 79.84±1.81% 48.79±4.41%
VTN-SlowFast-ViT 0.12±0.01 0.5689±0.0555 80.30±2.26% 46.47±6.91%
EPIC-Tent TSN [58] 0.17±0.01 0.2307±0.0505 54.04±6.03% 28.41±5.72%
SlowFast-R50 [16] 0.14±0.01 0.2887±0.0721 57.07±10.08% 31.09±4.98%
VTN-SlowFast-ViT 0.13±0.01 0.3318±0.0409 63.29±1.23% 34.95±3.17%
EPIC-Tower TSN [58] 0.24±0.05 0.4168±0.0616 71.06±2.54% 37.06±2.08%
SlowFast-R50 [16] 0.15±0.01 0.4936±0.0430 74.86±2.73% 40.44±10.04%
VTN-SlowFast-ViT 0.13±0.01 0.5689±0.0509 77.58±0.47% 50.20±6.50%
Table 20: Experiment results using one frame as input. The original experiment results for the classification task compared with the one-frame experiment results of training and testing the models using only one frame. We report the results by calculating the mean and standard deviation (StdDev) over the four splits. $\Delta$ Mean = Mean(train&test) - Mean(test only).
Activities Models Top1 Accuracy Rate (%)
Binary Classification Four-way Classification
Only Test Train & Test $\Delta$ Mean Only Test Train & Test $\Delta$ Mean
StdDev Mean StdDev Mean StdDev Mean StdDev Mean
Pipes-Struggle TSN [58] 2.10% 74.54% 2.10% 74.54% -3.57% 48.59% 3.57% 48.59% -%
One-frame test 25% 2.99% 64.60% 2.65% 67.18% 2.58% 3.87% 38.55% 3.00% 42.10% 3.55%
One-frame test 50% 2.80% 67.10% 3.35% 69.07% 1.97% 2.95% 40.85% 3.57% 42.31% 1.46%
One-frame test 75% 3.79% 67.39% 4.63% 70.19% 2.81% 3.49% 41.73% 3.58% 42.33% 0.61%
SlowFast-R50 [16] 2.44% 78.01% 2.44% 78.01% -1.97% 50.77% 1.97% 50.77% -%
One-frame test 25% 5.38% 63.24% 3.03% 63.25% 0.01% 3.90% 38.58% 1.85% 39.51% 0.94%
One-frame test 50% 4.39% 62.25% 3.11% 63.66% 1.41% 3.23% 38.67% 2.02% 39.79% 1.11%
One-frame test 75% 7.45% 62.31% 2.95% 63.36% 1.05% 3.55% 40.36% 1.89% 39.60% -0.75%
SlowFast-gMLP 1.62% 78.59% 1.62% 78.59% -2.97% 49.70% 2.97% 49.70% -%
One-frame test 25% 3.72% 62.60% 3.24% 64.92% 2.32% 6.99% 31.34% 2.14% 38.84% 7.50%
One-frame test 50% 3.72% 62.60% 3.68% 67.74% 5.14% 7.46% 31.73% 2.69% 41.87% 10.15%
One-frame test 75% 3.72% 62.60% 2.40% 65.24% 2.64% 6.40% 29.62% 1.99% 40.74% 11.12%
VTN-SlowFast-ViT 1.66% 77.80% 1.66% 77.80% -1.58% 50.83% 1.58% 50.83% -%
One-frame test 25% 3.83% 62.89% 3.27% 63.98% 1.09% 4.26% 34.10% 1.97% 39.82% 5.72%
One-frame test 50% 3.73% 62.79% 3.04% 63.77% 0.98% 4.58% 35.36% 2.42% 39.69% 4.33%
One-frame test 75% 3.73% 62.21% 2.75% 63.66% 1.45% 4.87% 34.14% 1.84% 40.24% 6.10%
Tent-Struggle TSN [58] 5.75% 68.05% 5.75% 68.05% -3.87% 38.23% 3.87% 38.23% -%
One-frame test 25% 6.09% 58.49% 3.06% 59.58% 1.10% 5.46% 31.77% 4.48% 37.86% 6.09%
One-frame test 50% 3.74% 61.88% 6.34% 61.07% -0.81% 5.07% 31.66% 4.38% 36.81% 5.15%
One-frame test 75% 3.33% 58.72% 5.32% 61.84% 3.12% 5.50% 33.59% 4.07% 35.79% 2.20%
SlowFast-R50 [16] 5.98% 69.56% 5.98% 69.56% 0.00% 4.68% 39.48% 4.68% 39.48% -%
One-frame test 25% 3.08% 55.82% 2.64% 55.24% -0.57% 3.28% 29.67% 4.73% 38.02% 8.35%
One-frame test 50% 2.69% 55.79% 4.03% 58.22% 2.43% 5.64% 29.20% 3.82% 36.32% 7.12%
One-frame test 75% 2.43% 55.07% 4.00% 56.82% 1.75% 4.24% 29.33% 4.59% 35.27% 5.93%
SlowFast-gMLP 6.16% 68.00% 6.16% 68.00% -1.96% 39.52% 1.96% 39.52% -%
One-frame test 25% 3.37% 52.38% 2.99% 55.64% 3.26% 8.21% 31.66% 3.07% 39.04% 7.38%
One-frame test 50% 7.84% 52.97% 4.63% 57.53% 4.57% 5.59% 33.11% 4.21% 34.48% 1.37%
One-frame test 75% 2.99% 53.36% 5.69% 52.94% -0.42% 5.87% 32.47% 2.89% 33.00% 0.53%
VTN-SlowFast-ViT 5.70% 68.53% 5.70% 68.53% -3.56% 40.11% 3.56% 40.11% -%
One-frame test 25% 2.84% 54.76% 3.73% 56.53% 1.77% 6.31% 33.06% 4.51% 37.16% 4.10%
One-frame test 50% 3.03% 55.84% 2.71% 55.27% -0.56% 7.26% 33.45% 4.07% 35.47% 2.02%
One-frame test 75% 3.07% 55.13% 3.50% 54.34% -0.79% 3.32% 35.69% 3.79% 35.95% 0.25%
Tower-Struggle TSN [58] 4.41% 82.59% 4.41% 82.59% -6.72% 59.49% 6.72% 59.49% -%
One-frame test 25% 6.40% 77.05% 3.39% 79.74% 2.69% 4.85% 54.50% 4.73% 57.56% 3.07%
One-frame test 50% 5.85% 73.66% 3.39% 79.74% 6.07% 7.02% 51.95% 4.73% 57.56% 5.61%
One-frame test 75% 3.65% 72.64% 3.39% 79.74% 7.10% 6.31% 51.50% 4.73% 57.56% 6.07%
SlowFast-R50 [16] 3.36% 88.24% 3.36% 88.24% -1.91% 63.92% 1.91% 63.92% -%
One-frame test 25% 3.45% 77.61% 4.28% 81.89% 4.28% 6.29% 55.34% 4.73% 57.56% 2.23%
One-frame test 50% 4.29% 80.48% 4.21% 81.98% 1.49% 4.43% 55.54% 4.73% 57.56% 2.02%
One-frame test 75% 3.01% 78.99% 3.25% 80.74% 1.75% 2.73% 55.78% 4.73% 57.56% 1.79%
SlowFast-gMLP 2.75% 87.86% 2.75% 87.86% -2.57% 66.40% 2.57% 66.40% -%
One-frame test 25% 3.39% 79.74% 3.39% 79.74% 0.00% 4.73% 57.56% 4.60% 58.44% 0.87%
One-frame test 50% 3.39% 79.74% 3.39% 79.74% 0.00% 4.73% 57.56% 4.00% 58.06% 0.50%
One-frame test 75% 3.39% 79.74% 3.39% 79.74% 0.00% 4.73% 57.56% 5.23% 58.81% 1.25%
VTN-SlowFast-ViT 4.54% 88.11% 4.54% 88.11% -4.09% 64.03% 4.09% 64.03% -%
One-frame test 25% 2.94% 79.35% 3.39% 79.74% 0.38% 4.73% 57.56% 4.73% 57.56% 0.00%
One-frame test 50% 2.94% 79.35% 3.39% 79.74% 0.38% 4.73% 57.56% 4.73% 57.56% 0.00%
One-frame test 75% 3.39% 79.74% 3.39% 79.74% 0.00% 4.73% 57.56% 4.73% 57.56% 0.00%
Table 21: Experiment results using shuffled frames. Mean Absolute Error (MAE) and Spearman’s Rank Correlation (Spearman’s Rho) with the corresponding converted classification accuracy for binary classification and four-way classification.
Activities Models Top1 Accuracy Rate (%)
Binary Classification Four-way Classification
Split 1 Split 2 Split 3 Split 4 StdDev Mean Split 1 Split 2 Split 3 Split 4 StdDev Mean
Pipes-Struggle SlowFast-R50 [16] 77.09% 75.11% 80.74% 79.09% 2.44% 78.01% 53.46% 50.66% 48.77% 50.19% 1.97% 50.77%
Shuffle frames test 64.22% 46.46% 60.25% 47.53% 8.96% 54.62% 24.36% 27.60% 30.08% 28.52% 2.41% 27.64%
Shuffle Train&Test 75.64% 74.67% 76.23% 78.33% 1.55% 76.22% 48.36% 42.36% 42.62% 49.05% 3.60% 45.60%
SlowFast-gMLP 78.91% 76.42% 80.33% 78.71% 1.62% 78.59% 53.46% 48.04% 46.72% 50.57% 2.97% 49.70%
Shuffle frames test 41.24% 45.15% 62.62% 47.30% 9.37% 49.08% 28.51% 22.97% 29.67% 22.66% 3.66% 25.95%
Shuffle Train&Test 74.91% 74.24% 75.92% 74.90% 0.69% 74.99% 47.27% 44.98% 46.72% 50.95% 2.51% 47.48%
VTN-SlowFast-ViT 77.82% 75.55% 79.51% 78.33% 1.66% 77.80% 50.54% 51.53% 48.77% 52.47% 1.58% 50.83%
Shuffle frames test 45.09% 53.97% 63.11% 43.42% 9.08% 51.40% 23.27% 32.40% 32.30% 24.18% 4.99% 28.04%
Shuffle Train&Test 73.46% 70.74% 73.77% 78.33% 3.15% 74.08% 46.18% 33.19% 40.98% 44.49% 5.77% 41.21%
Tent-Struggle SlowFast-R50 [16] 76.81% 72.11% 64.38% 64.94% 5.98% 69.56% 45.65% 38.10% 39.73% 34.42% 4.68% 39.48%
Shuffle frames test 50.00% 42.99% 43.56% 48.70% 3.55% 46.31% 22.32% 19.86% 22.60% 26.23% 2.62% 22.75%
Shuffle Train&Test 73.19% 75.51% 65.75% 57.14% 8.29% 67.90% 39.13% 39.46% 36.30% 31.17% 3.84% 36.52%
SlowFast-gMLP 73.19% 73.47% 63.01% 62.34% 6.16% 68.00% 41.30% 40.82% 36.99% 38.96% 1.96% 39.52%
Shuffle frames test 51.16% 45.03% 45.48% 49.74% 3.06% 47.85% 27.97% 29.25% 21.92% 24.42% 3.34% 25.89%
Shuffle Train&Test 68.84% 74.15% 65.07% 61.69% 5.34% 67.44% 36.96% 36.74% 31.51% 33.12% 2.70% 34.58%
VTN-SlowFast-ViT 73.91% 72.79% 65.07% 62.34% 5.70% 68.53% 42.75% 40.14% 42.47% 35.06% 3.56% 40.11%
Shuffle frames test 51.88% 42.99% 44.52% 48.70% 4.04% 47.02% 28.70% 29.93% 25.89% 24.68% 2.43% 27.30%
Shuffle Train&Test 68.84% 72.79% 63.01% 59.09% 6.08% 65.93% 41.30% 38.10% 30.14% 28.57% 6.15% 34.53%
Tower-Struggle SlowFast-R50 [16] 89.23% 85.18% 92.54% 86.00% 3.36% 88.24% 66.15% 64.82% 62.69% 62.00% 1.91% 63.92%
Shuffle frames test 32.00% 39.63% 49.25% 72.40% 17.54% 48.32% 21.54% 21.48% 42.99% 27.20% 10.15% 28.30%
Shuffle Train&Test 81.54% 77.78% 83.58% 82.00% 2.46% 81.23% 63.08% 50.00% 55.22% 62.00% 6.13% 57.58%
SlowFast-gMLP 89.23% 85.18% 91.04% 86.00% 2.75% 87.86% 64.62% 64.82% 70.15% 66.00% 2.57% 66.40%
Shuffle frames test 47.08% 42.96% 64.18% 64.00% 11.14% 54.56% 28.92% 22.59% 46.57% 36.40% 10.31% 33.62%
Shuffle Train&Test 81.54% 75.93% 82.09% 66.00% 7.46% 76.39% 63.08% 48.15% 65.67% 44.00% 10.75% 55.23%
VTN-SlowFast-ViT 89.23% 85.18% 94.03% 84.00% 4.54% 88.11% 67.69% 59.26% 67.16% 62.00% 4.09% 64.03%
Shuffle frames test 49.54% 42.22% 62.69% 66.40% 11.29% 55.21% 26.46% 19.63% 40.60% 38.00% 9.85% 31.17%
Shuffle Train&Test 87.69% 77.78% 86.57% 84.00% 4.43% 84.01% 60.00% 55.56% 62.69% 58.00% 3.02% 59.06%